OpenStudy (anonymous):

3x + 4y = -2 and 2x - 3y = 10

OpenStudy (anonymous):

What do you need?

OpenStudy (anonymous):

i just need the number for x and y but they both have to be the same for both equations

OpenStudy (anonymous):

3x-4y=8 multiply by 3 so 9x-12y=24 2x+3y=9 multiply by 4 so 8x+12y=36 add them 17x=60 x = 60/17 120/17 +3y=9 y=11/17